1) What is E0 at a point 2.40m from a light source whose output is 18 W? Assume the bulb emits radiation of a single frequency uniformly in all directions. (include appropriate units)


2) What is B0 at a point 2.40m from the light source? (include appropriate units)
Answer #1

Intensity I = power/area = Emax^2/2uoC

so Emax^2 = power* 2uoC/area

Emax^2 = 18 * 2* 4pi *10^-7 * 3e8/(pi * 2.4*2.4)

Emax^2 =750.21

Emax = 27.39 N/C


b.B0 = E/c = 27.39/3e8 = 9.13*10^-8 T